TERRY R. GENSYM

Work Address MIT Room E34–660 77 Massachusetts Ave. Cambridge, MA 02139

(617) 253–7955 Home Address 72 West St. Apt. 1 Cambridge, MA 02138 (617) 492–5543

Objective A large office, good pay, and very little work. Frequent expense-account trips to exotic lands would be a plus.

Experience CHENG’S CHINESE RESTAURANT Cambridge, MA 1970–1979 Soup taster for a small family-owned restaurant. Had final responsibility for the amount of Soy Sauce that went into the soups.

CARL’S FASHIONS Watertown, MA 1979–1982 Needle-threader for Kelly Hormel, a leading tailor in the Boston area. Duties included keeping all needles ready for use and threaded with the appropriate type and color of thread for each day’s jobs.

INSIDIOUS BLUE MACHINES San Jose, CA 1982–present Chief light bulb changer, with over 1500 square feet of office under my jurisdiction. The eyes of thousands were on me.

Education  MASSACHUSETTS INSTITUTE OF TECHNOLOGY  Cambridge, MA

B.A. in English History June 1970. Graduated at the top of my class.

Publications “A Multi-Threaded System for Needle Management,” Womens Wear Daily, August 20, 1981.

“The Representation of Flavor,” with August Rancatori, IEEE 1985 Proceedings of the Workshop on Taste in Computers.